# Bounds on the photon mass via the Shapiro effect in the solar system

### v2

@@ II.1 Deflection angle
Equation (10) matches the results from Refs. Accioly1 ; Accioly2 calculated via scattering methods, as well as those from a simplified calculation using a straight-line approximation, cf. App. A. In Ref. Accioly2 the authors use 43-GHz data from the Very-Long Baseline Array (VLBA) Fomalont and derive $m_{\gamma}\lesssim 3.5\times 10^{-5}\,{\rm eV/c^{2}}$, whereas in Ref. Accioly1 interferometric techniques are also used, but with 2-GHz radio waves Lebach , to obtain [^3] In Ref. Accioly1 the authors quote $m_{\gamma}\lesssim 10^{-40}\,{\rm g}\approx 5.6\times 10^{-8}\,{\rm eV/c^{2}}$. However, the deviation of the measured angles relative to Einstein’s result is $\Delta/\delta_{\rm E}\approx 8\times 10^{-4}$ Lebach , so that using the first equation in their section 3 we get $m_{\gamma}\lesssim 5.9\times 10^{-40}\,{\rm g}\approx 3.3\times 10^{-7}\,{\rm eV/c^{2}}$, a factor six worse than their stated limit. $m_{\gamma}\lesssim 3.3\times 10^{-7}\,{\rm eV/c^{2}}$.
@@ III Bound on the photon mass via Cassini data
The analysis of the time-delay data performed in Ref. Cassini_2 was based on the accurate characterization of Cassini’s orbital dynamics. This was done using the Orbit Determination Program from NASA’s Jet Propulsion Laboratory with which a fit of the data with up to 12 free parameters, including the initial state vector of the Cassini spacecraft, non-gravitational accelerations, thermo-optical properties of the antenna and Eddington’s $\gamma$ parameter, was performed. The orbital parameters were determined with uncertainties in position $\lesssim 50$ km for geocentric distances of $\approx 8.4$ AU. Furthermore, the uncertainty in the FFS, given by the Allan deviation, was found to be $\sigma_{y}\approx 2\times 10^{-14}$, practically constant and a factor $\approx 10^{4}$ smaller than the measured and predicted FFS Cassini_1 ; Cassini_2 ; Cassini_imp_param ; Cassini_error_budget .

### v4

@@ II.1 Deflection angle
Equation (10) matches the results from Refs. Accioly1 ; Accioly2 calculated via scattering methods, as well as those from a simplified calculation using a straight-line approximation, cf. App. C. In Ref. Accioly2 the authors use 43-GHz data from the Very-Long Baseline Array (VLBA) Fomalont and derive $m_{\gamma}\lesssim 3.5\times 10^{-5}\,{\rm eV/c^{2}}$, whereas in Ref. Accioly1 interferometric techniques are also used, but with 2-GHz radio waves Lebach , to obtain [^3] In Ref. Accioly1 the authors quote $m_{\gamma}\lesssim 10^{-40}\,{\rm g}\approx 5.6\times 10^{-8}\,{\rm eV/c^{2}}$. However, the deviation of the measured angles relative to Einstein’s result is $\Delta/\delta_{\rm E}\approx 8\times 10^{-4}$ Lebach , so that using the first equation in their section 3 we get $m_{\gamma}\lesssim 5.9\times 10^{-40}\,{\rm g}\approx 3.3\times 10^{-7}\,{\rm eV/c^{2}}$, a factor six worse than their stated limit. $m_{\gamma}\lesssim 3.3\times 10^{-7}\,{\rm eV/c^{2}}$.
@@ III Bound on the photon mass via Cassini data
The analysis of the time-delay data performed in Ref. Cassini_2 was based on the accurate characterization of Cassini’s orbital dynamics. This was done using the Orbit Determination Program from NASA’s Jet Propulsion Laboratory with which a fit of the data with up to 12 free parameters, including the initial state vector of the Cassini spacecraft, non-gravitational accelerations, thermo-optical properties of the antenna and Eddington’s $\gamma$ parameter, was performed (see App. D). The orbital parameters were determined with uncertainties in position $\lesssim 50$ km for geocentric distances of $\approx 8.4$ AU. Furthermore, the uncertainty in the FFS, given by the Allan deviation, was found to be $\sigma_{y}\approx 2\times 10^{-14}$, practically constant and a factor $\approx 10^{4}$ smaller than the measured and predicted FFS Cassini_1 ; Cassini_2 ; Cassini_imp_param ; Cassini_error_budget .
@@ Appendix C Deflection angle in the straight-line approximation
Appendix D Impact of a photon mass on the orbit determination of the Cassini spacecraft and parameter fit
The mass of the photon, $m_{\gamma}$, is small enough to make the photon a test particle in the solar gravitational background: the spacetime metric is independent of $m_{\gamma}$. The consequence is that the geodesics remain independent of $m_{\gamma}$. Therefore, on the theoretical side, the determination of the orbital parameters by means of solving the equations of motion is unaffected by $m_{\gamma}$. We are nonetheless dealing with a spacecraft moving at interplanetary distances whose trajectory must be precisely controlled to keep it on track.
The communication between spacecraft and ground station is performed using electromagnetic signals. It is thus plausible that a finite photon mass will have some impact on the measurement and subsequent reconstruction of the orbital parameters. As will be shown, the range and range rate are influenced by a finite $m_{\gamma}$, since these depend on the two-way light time. Let us first briefly address the orbit determination procedure.
Orbit determination is an iterative process starting with the nominal orbit involving the simultaneous integration of the equations of motion of all relevant participants (including all known influences) Moyer ; Thornton . The determination of the trajectory is performed by comparing observed quantities, obtained via tracking (range and range rate) to calculated ones. Using suitable fitting procedures the optimal model parameter set is found and the residuals are minimized. The parameters estimated in this way are fed back into the model and the process is repeated until adequate convergence is achieved.
The model parameters (a total of 12) that are estimated in the fit are Cassini_1 ; Cassini_2 : the six components of the spacecraft state vector at a reference epoch, the three components of the induced acceleration from the radioisotope thermal generators, the specular and diffuse reflectivity of the spacecraft high gain antenna, and the relativistic PPN parameter $\gamma$. If we were to repeat the orbital determination procedure, we would fix $\gamma=1$ and include the photon mass as a free parameter.
Range and range rate (Doppler) measurements inform about distance and velocity relative to the tracking station on Earth. The spacecraft range is measured using the transit time of a ranging signal, $\Delta\tau$, which is sent from a ground station, electronically processed on-board and re-emitted to Earth. The range, $\rho$, is given by $\rho=c\Delta\tau$, whereas the range rate, $\dot{\rho}$, is determined via Doppler tracking: $f_{\rm R}=f_{\rm T}\left(1\pm\dot{\rho}/c\right)$, where $f_{\rm T}$ and $f_{\rm R}$ are the transmitted and received frequencies, respectively (the plus or minus sign depends on the sign of the relative range rate). Due to the dependence on the travel time (one way or two way), the modelled distances and velocities entering the orbital determination fit are potentially affected by a finite photon mass. This issue will be discussed below.
Frequencies are measured by comparison with hydrogen masers providing an extremely stable time standard. Even if a finite photon mass would have sensible effects on the hydrogen spectrum, the sent and received frequencies are both measured against the same reference frequency and the Doppler shifts are not large enough to produce meaningful effects. Therefore, the photon mass does not play a role in frequency comparison.
The re-transmission of the radio signal is performed as follows. The signal received at the spacecraft is already Doppler shifted (one-way Doppler shift). The transponder on-board, the Ka band translator, KaT, is responsible for collecting the incoming radio signal, measure its frequency, lock it in and then re-emit a signal with this one-way Doppler-shifted frequency back to Earth. The signal received at the ground station is therefore the two-way Doppler shifted signal, which contains the expected gravitationally delayed frequency, but also several other effects that must be removed, such as non-gravitational accelerations, coronal plasma effects and tropospheric effects. Despite the fact that some of these are measured using electromagnetic radiation (atmospheric measurements use radar), the very short times and distances involved lead us to not expect any sensitivity to a finite photon mass. Therefore, also in the re-transmission process the photon mass will have no impact.
Regarding non-gravitational accelerations, the only forces with potential impact on the experiment were the solar radiation pressure and the thermal emission from the radioisotope thermal generators Cassini_error_budget . The radiation pressure, which would have been affected by a photon mass, induces a non-gravitational acceleration a hundred times smaller than that due to thermal emissions, which were themselves found to be very small. These two effects are overall negligible and any correction from a finite photon mass would not change that.
We have assumed that a finite mass would have no meaningful impact in the orbital parameters, thereby justifying the usage of the Cassini data as presented. This approach would be strictly rigorous if the effects of a finite photon mass would be smaller than the errors in the determination of the orbital parameters, such as heliocentric or geocentric distance of the Cassini spacecraft.
The current best limit is $m_{\gamma}<10^{-18}\,{\rm eV/c^{2}}$ PDG ; Ryutov , which gives $\alpha^{2}<10^{-28}$, cf. eq. (5) with $f_{\rm T}=8.4$ GHz. From $v=c\sqrt{1-\alpha^{2}}$ we may write $v=c-\Delta c$ with $\Delta c=c\alpha^{2}/2$ and find
$\frac{c-v}{c}=\frac{\Delta c}{c}\approx 10^{-28}\,.$ (50)
Measuring this fractional difference is impossible with current or near future capabilities.
The orbital distances have been determined with unprecedented accuracy. The spacecraft position was determined with errors $\Delta L\approx 50$ km at geocentric distances of $L\approx 8$ AU MORE ,. We thus have
$\frac{L-(L\pm\Delta L)}{L}=\pm\frac{\Delta L}{L}\approx\pm 10^{-8}\,.$ (51)
Despite being an impressive feat, this is several orders of magnitude larger than the order of magnitude obtained in eq. (50).
How large would $m_{\gamma}$ have to be to produce a similar relative variation? Assuming that range measurements (or determinations via fits) would be affected by a finite photon mass, we may approximately write $\Delta L/L\approx\alpha^{2}$. Therefore, if the uncertainty in the position determination remains roughly the same, we must have $\alpha^{2}<10^{-8}$, which leads to $m_{\gamma}<10^{-9}\,{\rm eV/c^{2}}$, roughly hundred times smaller than our upper bound. One could thus say that, had we used the full orbital determination program including all the dynamical variables used in the Cassini study, but now with the extra free parameter being $m_{\gamma}$ instead of $\gamma$, the fit could have (optimistically) returned an upper bound at this ballpark. Note that this takes only the position uncertainty into account, but there might be other larger sources of uncertainty, which would cause this limit to degrade further.
In conclusion, our bound is a conservative estimate. A more thorough – but technically out of scope – treatment could lead to a better limit by a factor $\sim 100$, probably less. As shown above, the heliocentric distances and other orbital parameters extracted from NASA’s Horizon calculator and from the Cassini solar conjunction experiment are not expected to significantly change (within the stated relative uncertainties in range) if $m_{\gamma}<10^{-9}\,{\rm eV/c^{2}}$. Our approach does not consider these tighter relative uncertainties, thus leading to the upper bound $m_{\gamma}<4.9\times 10^{-7}\,{\rm eV/c^{2}}$, representing a conservative limit that is compatible with the larger relative uncertainties in the FFS after the nominal fit was performed. This nonetheless encompasses the estimate given above. Therefore, we believe that there is a chance that our bound could improve by at most a couple of orders of magnitude if a re-fit would be feasible.